Use inductive reasoning to predict the next number in each sequence.
3, 7, 11, 15, _____________
1, -2, 4, -8 ________________

Found 4 solutions by stanbon, yannick, scott8148, Edwin McCravy:
Answer by stanbon(75887)

predict the next number in each sequence.
3, 7, 11, 15, 15+4=19
1, -2, 4, -8 , (-2)^4 = 16
====================
Cheers,
stan H.

Answer by yannick(6)

3, 7, 11, 15, _____________
The difference between the numbers is exactly 4.
3+4=7, 7+4=11, 11+4=15, 15+4=19, 19+4=23, etc___________________


1, -2, 4, -8 ________________
1-3=-2, -2 +(-3*-2=-6)=4, 4+(-6*2=-12)=-8, -8+(-12*-2=24)=16, etc...

Answer by scott8148(6628)

in the first sequence, you add 4 to get from one term to the next ... so the missing term is 19

in the second sequence, you multiply by -2 to get the next term ... so the missing term is 16

Answer by Edwin McCravy(20086)


Use inductive reasoning to predict the next 
number in each sequence. 
3, 7, 11, 15, _____________ 

To get from 3 to 7 you either have to 

1. Add 4
or
2. Multiply by 7/3

To get from 7 to 11 you either have to 

1. Add 4
or
2. Multiply by 11/7

To get from 11 to 15 you either have to 

1. Add 4
or
2. Multiply by 15/11

The only rule that was the same on all those 
was "Add 4".

So we add 4 to 15 and get 19.

Answer

3, 7, 11, 15, _19_

==============================================

1, -2, 4, -8 ________________

To get from 1 to -2 you either have to 

1. Add -3
or
2. Multiply by -2

To get from -2 to 4 you either have to 

1. Add 6
or
2. Multiply by -2

To get from 4 to -8 you either have to 

1. Add -12
or
2. Multiply by -2

The only rule that was the same on all those 
was "Multiply by -2".

So we multiply -8 by -2 and get 16.

Answer:

1, -2, 4, -8, _16_

Edwin